### v1

# RoundaboutHD: High-Resolution Real-World Urban Environment Benchmark for Multi-Camera Vehicle Tracking

Abstract
The multi-camera vehicle tracking (MCVT) framework holds significant potential for smart city applications, including anomaly detection, traffic density estimation, and suspect vehicle tracking. However, current publicly available datasets exhibit limitations, such as overly simplistic scenarios, low-resolution footage, and insufficiently diverse conditions, creating a considerable gap between academic research and real-world scenario. To fill this gap, we introduce RoundaboutHD, a comprehensive, high-resolution multi-camera vehicle tracking benchmark dataset specifically designed to represent real-world roundabout scenarios. RoundaboutHD provides a total of 40 minutes of labelled video footage captured by four non-overlapping, high-resolution (4K resolution, 15 fps) cameras. In total, 512 unique vehicle identities are annotated across different camera views, offering rich cross-camera association data. RoundaboutHD offers temporal consistency video footage and enhanced challenges, including increased occlusions and nonlinear movement inside the roundabout. In addition to the full MCVT dataset, several subsets are also available for object detection, single camera tracking, and image-based vehicle re-identification (ReID) tasks. Vehicle model information and camera modelling/ geometry information are also included to support further analysis. We provide baseline results for vehicle detection, single-camera tracking, image-based vehicle re-identification, and multi-camera tracking. The dataset and the evaluation code are publicly available at:https://github.com/siri-rouser/RoundaboutHD.git
1 Introduction
With the continuous expansion of urbanization, the city traffic management is becoming more and more challenging, which is calling for the development of intelligent transportation systems (ITS) within the broader concept of smart cities. A practical approach to construct real-world ITS is leveraging existing traffic cameras, which capture extensive amounts of useful information in their video footage.
Figure 1: MCVT Problem Overview. The example shows two tracklet from two different non-overlapping FOV camera (Cam01 and Cam02). The cropped images above each camera represent tracklets generated from detection and single-camera tracking, and the blue arrow represents the cross-camera vehicle ReID process, linking the two tracklets as the same vehicle identity across views.
Dataset Cam_Num Boxes Average Duration Resolution Fps Video Vehicle Info Open-Sourced
Image- based VehicleID [20] / 221,763 / / / $\times$× $\surd square-root\surd$
[…]
Vehicle-1M [12] / 936,051 / / $\times$× $\surd square-root\surd$
[…]
SyntheVehicle [15] 340 4,623,184 3min 1920×1080 10 $\surd ×\times× square-root\surd$
LHTV [35] 16 / 63.26min 1920×1080 30 $\surd ×\times× square-root\surd$
HST [38] 6 / 30.33min / 30 $\surd ×\times× square-root\surd$
[…]
A common application of these video resources is tracking vehicles across multiple camera views, which provides vehicle trajectory information to city management authorities. This task, known as multi-target multi-camera tracking (MTMCT), has garnered considerable research interest in recent years [2]. Generally, this MTMCT problem can be divided into 3 distinct sub-problems, as shown in Fig. 1. (1) Object Detection, which detects and identifies vehicles within the camera scenes; (2) Single-Camera Tracking, which tracks vehicles’ movement across sequential frames within each camera’s field of view; and (3) Multi-Camera Re-identification, which matches and associate’s vehicle identities across separate different camera views. This identification can be performed either at an image-level (image-based re-identification) or at a tracklet-level (video-based re-identification), ultimately reconstructing complete vehicle trajectories.
Each of these sub-problems presents significant challenges in its respective domain, leading to the creation of specialized datasets for object detection, tracking, and re-identification tasks [18, 7, 22]. However, there only have limited number of open-source real-world vehicle-specified MTMCT dataset. To the best of the authors’ knowledge, CityFlow [31] is currently the only open-source dataset of this kind.
The main challenges for making a real-world MCVT datasets are the small inter-class variability and large intra-class variability. Vehicles of the same model typically exhibit very subtle differences, making the re-identification process very challenging even for humans. Conversely, the same vehicle can have dramatic appearance difference due to different lighting condition, camera viewpoint and occlusion, which further complicating the match process. Moreover, the dataset creation is also very time-consuming, requiring the annotation and label the link of hundreds or even thousands of vehicles over extended periods. Considering these challenges, we introduce RoundaboutHD to enrich the available MCVT datasets, the geographic layout is illustrated in Fig. 2. RoundaboutHD comprises 40 minutes of annotated video footage recorded simultaneously from four non-overlapping high-resolution cameras (4K, 15 FPS), with 10 minutes of labeled video per camera. The dataset also introduces increased challenges, including dramatic appearance change due the change of vehicle view angle, frequent occlusions and nonlinear vehicle movements. In addition to the primary dataset designed for evaluating MCVT, we provide subsets for object detection, single camera tracking and image-based ReID. Detailed vehicle attributes (including vehicle type, color, make, and model) with comprehensive camera geometry/calibration information are also provided, enabling diverse research within this domain.
Compared to the widely used CityFlow benchmark, RoundaboutHD focus on a single, continuous scene to ensure high temporal consistency, while maintaining a comparable number of annotated vehicle identities (512 vs. 666). Additionally, it emphasizes real-world representativeness by offering longer, higher-resolution video, and richer contextual annotations—including vehicle attributes and camera calibration information. We hope that the release of this dataset will support and push forward research in ITS, especially for the real-world application of MCVT. The dataset, along with ground-truth annotations for detection, single-camera tracking, image-based ReID, and multi-camera tracking, as well as the associated evaluation tools, are publicly available.
2 Related Work
This section reviews and compares several widely used benchmarks for the ReID task, as summarized in Tab. 1. The table splits existing ReID datasets into image-based and video-based (MTMCT), covering vehicle-specified re-identification scenarios.
2.1 Vehicle Image-Based ReID Dataset
The vehicle images-based ReID datasets [22, 23, 21, 20, 16, 12] are designed specifically for ReID tasks at image level, excluding detection and tracking stages. Among those datasets, VeRI-776 [21] is one of the most commonly used datasets for image-based vehicle ReID. It is an extended version of the original VeRI dataset [22], containing images of 776 vehicles captured by different cameras. VERI-WILD [23] is notable for its large scale, comprising images collected from 174 surveillance cameras over one month in uncontrolled real-world conditions. PKU-VehicleID [20] stands out for its inclusion of detailed vehicle model annotations, offering richer contextual information than many other datasets. The VRIC [16] contributes through its realistic and unconstrained variations in resolution, with all images retrieved from the UA-DETRAC vehicle tracking dataset [34].Vehicle-1M [12] is the largest vehicle ReID dataset published to date, containing nearly 1 million images of 55,527 vehicles across 400 different models. These datasets provide a solid foundation for vehicle ReID research and have helped drive progress toward video-based and multi-camera tracking methods.
2.2 Vehicle Video-Based ReID Dataset
Figure 2: Camera Layout and Preview of the Proposed RoundaboutHD Dataset. Distances among camera pairs are shown, and red arrows indicate the orientation of each camera
Compared to the image-based ReID task, the video-based ReID task adds the temporal consistency that makes it suitable for MTMCT tasks. While several open-source MTMCT datasets exist for person ReID research (e.g. [28, 40]), to the best of the authors’ knowledge, only two publicly available datasets works on vehicle-specified MTMCT: CityFlow [31] and Synthehicle [15]. CityFlow, released through the CVPR AI City Challenge workshop [26], is the most widely used benchmark in MCVT research. It contains 3.58 hours of video from 46 cameras with 229,680 bounding boxes of 666 vehicle identities annotated. However, each camera covers only a short duration, and the video quality is relatively low (960p), which limits its effectiveness in representing real-world urban environments. Synthehicle, on the other hand, is the first synthetic dataset that eliminate the need for manual labelling. Compared to other datasets, it offers a large-scale material with rich annotations, including 2D bounding boxes, 3D positions, segmentation masks, and depth information. Nevertheless, all video content is generated in the CARLA [9], introducing a domain gap that reduces its applicability to real-world scenarios. In addition to these open-source datasets, several private video-based ReID datasets exist for MCVT evaluation, such as the Large-scale High-resolution Traffic Video (LHTV) dataset [35] and the Highway Surveillance Traffic (HST) dataset [38]. However, these datasets are not publicly available, which restricts their use in broader academic research and benchmarking.
3 Dataset Analysis
3.1 Dataset Overview
Figure 3: Illustration of the Mapping Process for Cam01. The top diagram demonstrates how camera transformation matrices are used to map pixel locations to real-world coordinates (latitude and longitude). The bottom image shows the satellite view of Cam01’s coverage area (left) and a sample video frame (right). Colored point pairs represent corresponding GPS coordinates and pixel positions, with each pair connected by a yellow dashed line.
The proposed RoundaboutHD dataset was collected in a small town in the United States. The geographic layout, inter-camera distances, and sample views from each camera are illustrated in Fig. 2. The maximum distance between two cameras is approximately 0.92 km. Several retail parks and side roads are between camera pairs, increasing the uncertainty of vehicle movement patterns and making the tracking environment closer to real-world urban traffic scenarios.
The total duration of the RoundaboutHD dataset is 40 minutes, each 4k 15fps camera contains 10 minutes videos. All videos begin recording simultaneously to support temporally aligned multi-camera tracking. Within the 40-minute recording, a total of 549,909 bounding boxes of 512 vehicles are manually annotated. The number of annotated vehicles per camera is summarized in Tab. 2.
cam01 cam02 cam03 cam04 Multi-Cam
Num of
Vehicle
260 242 278 255 310
Table 2: Number of Annotated Vehicles in Each Camera. Multi-Cam refers to vehicles appearing in at least two cameras.
Camera modelling parameters are provided based on the camera model in [10], through those parameters, user can transform point coordinates in the image to real-world or to geographic coordinates. An illustration of this mapping process is presented in Fig. 3. In addition, human annotators manually label vehicle attributes including colour, type, make, and specific model for all vehicles shown up in our dataset. This rich contextual information offers valuable support for a range of future research. Based on these annotations, the five most common vehicle makes are Chevrolet, Ford, Honda, Toyota, and Nissan, while the most frequently observed vehicle models are Chevrolet Silverado, Ford Transit, Honda CR-V, Chevrolet Equinox, and Toyota Camry. A statistical summary of vehicle types and colours is provided in Fig. 4.
Figure 4: The Distribution of Vehicle Colors and Types in terms of Vehicle Identities in RoundaboutHD. In the vehicle type chart, Car refers to small passenger vehicles, SUV denotes Sport Utility Vehicles, HGV represents Heavy Goods Vehicles, and MPV stands for Multi-Purpose Vehicles.
For privacy concerns, a specific angle of the cameras is pre-set to make sure no recognizable human face or readable license plate is recorded. A comparative overview of RoundaboutHD and existing datasets is provided in Tab. 1.
3.2 Dataset Annotation
Manual frame-level annotation for MCVT is highly time-consuming and labor-intensive. To improve efficiency, we adopt a semi-automatic annotation pipeline combining detection, feature extraction, and trajectory generation with manual refinement. First, the YOLOv12x [32] object detector pre-trained on COCO dataset [18] is employed to generate initial bounding boxes. A confidence threshold of 0.35 is applied, and detections with bounding box areas smaller than 900 pixels are filtered out to reduce noise from small or distant objects. All initial detection results are manually reviewed to correct false positives and false negative identities. Next, deep appearance features are extracted using a ResNet-101 [13] model. These features, along with the refined detections, are input to SMILEtrack [33] to get initial trajectories. The resulting trajectories are manually inspected, and ID switches are corrected to ensure accurate single-camera tracking. Based on above information, the cross-camera association is conducted manually. Vehicles are matched across different camera views based on a combination of visual attributes (colour, make, and model) and spatial-temporal consistency between camera pairs. A challenging cross-camera vehicle movement example is shown in Fig. 5. Finally, an image-based ReID subset is constructed based on the multi-camera association result. To ensure high-quality visual data, cropped images are extracted from bounding boxes with a minimum resolution of 3,500 pixels, sampled every 5 frames, to form the training and gallery sets. The query set is sampled every 45 frames with prioritizing high resolution cropped images. This approach ensures sufficient visual diversity and quality for robust vehicle ReID benchmarking.
Figure 5: Example of cross-camera vehicle trajectory annotation with timestamp information.
3.3 Subset Overview
During the construction of the main MCVT dataset, we generate three task-specific subsets to support a broad range of research applications: an object detection dataset, a single-camera tracking dataset, and an image-based vehicle ReID dataset.
Object detection dataset. This subset includes a total of 549,909 manually verified bounding boxes, each annotated with vehicle type labels. These annotations can be directly used for training vehicle-specific object detectors, vehicle type classifiers, and other vision-based traffic applications.
Single-camera tracking dataset. The SCT subset includes a total of 1,082 annotated trajectories, composed of 233,189 bounding boxes. While Table 2 provides the number of unique vehicle identities per camera, the total number of annotated trajectories slightly exceeds the number of unique identities. The difference is raised from occasional long-term occlusions, which leads to id switch within individual camera views. The distribution of annotated trajectories across the four cameras is summarized in Tab. 3:
cam01 cam02 cam03 cam04
Num of
Vehicle
262 244 279 257
Table 3: Number of Annotated Trajectories in Each Camera.
Image-based vehicle ReID dataset. We construct a high-quality image-based ReID subset that is comparable in both the number of images and unique identities to several widely used ReID benchmark datasets. The dataset consists of 65,528 cropped images in total. Among these, 23,227 gallery images and 1,173 query images from 310 unique identities form the testing set. The training set includes 41,128 images from 510 unique identities. On average, each identity is represented by approximately 128.5 images, providing rich appearance variation for robust ReID evaluation.
3.4 Evaluation
To facilitate standardized benchmarking, we provide an evaluation framework for the main dataset and all subsets as part of the package. This benchmark enables comparison of different algorithm by evaluating the result text file with ground truth using task-specific performance metrics. Details on the required input file formats for each task are available in our open-source GitHub repository.
Object detection evaluation. For object detection, we adopt standard evaluation metrics, including AP@0.5, AP@0.75, AP@0.9, and the mean Average Precision (mAP).
Single-camera tracking evaluation. For SCT, we select serval metrics used by the famous MOT Challenges [7], includes IDF1 score [29], ID Switches and MOTA [3] to evaluate the performance of each SCT algorithm.
Image-based vehicle ReID evaluation. For the ReID task, we follow the evaluation protocol used by the FastReID benchmark [14]. Performance is reported using three key metrics: the mAP, rank-k accuracy, mean Inverse Negative Penalty (mINP) [6].
Multi-camera vehicle tracking evaluation. For the main MCVT dataset, we adopt the evaluation format form the well-known CVPR AICITY challenge to ensure an easy and duplicatable comparison. The evaluation framework calculates the IDF1 score [29] (includes IDF1, IDP and IDR) as the metrics to reflect the algorithm performance.
Overall, this evaluation framework provides a unified, easy-to-use approach for benchmarking algorithms across all tasks on the RoundaboutHD dataset.
4 Experiment
In this section, we conducted several different experiments on object detection (Sec. 4.1), single camera tracking (Sec. 4.2), image-based ReID (Sec. 4.3), and multi-camera tracking (Sec. 4.4) using multiple SOTA methods to provide baseline performances for future research.
4.1 Object Detection
Object detection serves as a curial part in the whole MTMCT pipeline, as its quality directly impacts the accuracy of subsequent tracking and re-identification modules. The RoundaboutHD dataset presents a variety of challenging scenarios for object detection, including occlusion, varying lighting conditions, and uncommon vehicle types. Examples of these challenges are illustrated in Fig. 6.
Figure 6: Example Challenging Cases for Object Detection in RoundaboutHD. (a) A combination of rare vehicle types and severe occlusion leads to both false positives and false negatives. (b) Special vehicle appearances result in incomplete detections. (c) Heavy occlusion causes false negatives.
Object detectors are typically classified into two categories: one-stage detectors, such as YOLO [32], and two-stage detectors, such as Fast R-CNN [11]. Since 2020, one-stage detectors—particularly the YOLO family—have become dominant in both academic research and real-world applications due to their speed and accuracy. To evaluate how SOTA detectors perform under such conditions, we use the recently released attention-based YOLOv12x [32] model and transformer architecture based YOLOv11X [17] as a benchmark, setting confidence threshold to 0.3, non-maximum suppression (NMS) to be true, using an inference image size of 1280, and applying an IoU threshold of 0.4. Tab. 4 and Tab. 5 presents its performance on the RoundaboutHD dataset, offering insights into detection performance in complex urban scenes.
Camera AP@0.5 AP@0.75 AP@0.9 mAP
cam01 40.42 40.25 37.03 39.23
cam02 69.39 69.22 64.31 67.64
cam03 74.85 74.70 71.08 73.54
cam04 79.52 78.90 74.17 77.53
Table 4: YOLOv12x Object Detection Performance. All values are giving in percentage (%).
Camera AP@0.5 AP@0.75 AP@0.9 mAP
cam01 45.10 44.70 36.44 42.08
cam02 74.48 73.81 62.90 70.39
cam03 76.87 76.50 69.82 74.39
cam04 81.08 80.03 68.68 76.60
Table 5: YOLOv11x Object Detection Performance. All values are giving in percentage (%).
According to the results, the transformer-based YOLOv11 slightly outperforms YOLOv12, achieving a 1.38% higher average mAP. Notably, the ground truth annotations exclude static vehicles parked in the visible parking lot. However, the YOLO detector still identifies these vehicles, leading to a higher FP rate—particularly in Cam01, where multiple vehicles on the right side of the video footage are obvious. This affects the apparent detection performance and should be considered when interpreting the results.
Among the four cameras, Cam04 is the least affected by this issue and thus provides a more representative measure of detection accuracy. Incorporating a filtering mechanism to remove static vehicle detections should be considered for future experiments to obtain more reliable and meaningful evaluation results.
4.2 Single Camera Tracking
In the general MTMCT task, the performance of MTMCT largely depends on the performance of SCT, as errors in SCT can negatively affect downstream association across cameras. As a result, many researchers have focused on improving SCT to enhance the overall effectiveness of MTMCT systems. In our RoundaboutHD dataset, we provide some extra challenging for SCT. Due to the nature of roundabout, vehicles follow non-linear, curved trajectories rather than simple straight-line motion. In addition, a central statue located at the center of the roundabout (visible in Fig. 2) introduces frequent occlusions. These factors collectively increase the difficulty of tracking consistent vehicle identities over time within each camera view.
To evaluate SCT performance on RoundaboutHD, we follow the widely used tracking-by-detection paradigm provided by BOXMOT [4]. BOXMOT is an open-source pluggable SCT framework that enables easy and straightforward deployment of multiple different SOTA SCT algorithms with minimal configuration change. For consistency across evaluations, detection results from the YOLOv12x model are used as input for all trackers. This ensures that the evaluation focus on the performance differences of tracking algorithms themselves. Additionally, to avoid distractions from static vehicles parked in visible parking areas, trajectories corresponding to stationary objects are filtered out prior to evaluation. Detailed results from different trackers are listed from Tab. 6 to Tab. 10, and a performance summary is shown in Tab. 11.
Camera IDF1 IDP IDR IDs MOTA
cam01 62.1 86.0 48.6 19 43.9
cam02 66.5 92.7 51.8 11 49.4
cam03 41.9 85.1 27.8 7 24.0
cam04 69.0 89.3 56.2 29 54.0
Table 6: BoostTrack [30] Performance Evaluation.
Camera IDF1 IDP IDR IDs MOTA
cam01 73.9 75.2 72.6 127 80.1
cam02 92.2 95.3 85.7 9 83.3
cam03 78.5 80.6 76.5 94 83.7
cam04 86.1 89.4 83.0 36 83.4
Table 7: BotSort [1] Performance Evaluation.
Camera IDF1 IDP IDR IDs MOTA
cam01 71.8 75.0 68.9 127 77.4
cam02 87.6 95.2 81.1 11 80.1
cam03 76.4 79.0 73.9 112 81.1
cam04 85.7 90.3 81.6 43 83.2
Table 8: Deepocsort [25] Performance Evaluation.
Camera IDF1 IDP IDR IDs MOTA
cam01 72.6 72.2 73.1 129 79.0
cam02 90.5 93.3 87.8 10 83.7
cam03 84.2 86.9 81.7 39 82.4
cam04 86.0 86.8 85.2 45 85.3
Table 9: ByteTrack [39] Performance Evaluation.
Camera IDF1 IDP IDR IDs MOTA
cam01 70.3 81.1 62.0 112 71.2
cam02 79.3 96.7 67.3 14 68.0
cam03 75.8 85.6 68.1 93 76.4
cam04 79.9 91.6 70.8 43 74.8
Table 10: OCSort [5] Performance Evaluation.
Algorithm IDF1 IDP IDR IDs MOTA
BoostTrack 59.88 88.28 46.10 16.5 42.83
BotSort 82.68 85.13 79.45 66.5 82.63
DeepOCSort 80.38 84.88 76.38 73.25 80.45
ByteTrack 83.33 84.80 81.95 55.75 82.6
OCsort 76.33 88.75 67.05 65.5 72.6
Table 11: SCT Algorithms Performance Summary. All metrics in this table are the average values across four cameras.
Among all evaluated algorithms, ByteTrack achieves the highest average performance in terms of both IDF1 and MOTA. BoostTrack reports the lowest number of ID switches; however, this might because of its limited tracking performance, as indicated by its lower IDR score, which suggests fewer trajectories are counted in the result. For the comparison among different scene, trackers generally perform better in Cam02 and Cam04 compared to Cam01 and Cam03. This disparity is largely due to the presence of a large statue located at the center of the roundabout in Cam01 and Cam03, which blocks the view and introduce frequent occlusion during the tracking progress.
4.3 Image-Based ReID
The object re-identification is an image retrieval task and represents a key component of MTMCT. It aims to match a given query image to images in a gallery, identifying the most similar gallery image that corresponds to the same object identity. The primary difference between image-based ReID and video-based ReID is that the former associates identities represented by single images, while the latter associates identities represented by multiple sequential images (tracklets).
To evaluate the image-based vehicle ReID subset, we adopt the widely-used FastReID toolbox [14], which provides a straightforward framework for training and evaluating various SOTA algorithms. We select three popular baseline models commonly used for ReID task, BoT [24], AGW [37] and SBS (a FastReId improved baseline based on BoT) [14] for benchmarking the RoundaboutHD image-based ReID subset. To ensure a fair comparison, all three baselines utilize the same ResNet50-IBN backbone, pre-trained on ImageNet [8]. Each baseline model is trained on our training dataset using the combination of CrossEntropyLoss and TripletLoss with FastReID’s default training parameters for 60 epochs, and the best-performing model is selected for the final evaluation. All training is conducted on an NVIDIA GeForce RTX 4070 Ti GPU. The results are shown in Tab. 12 and Tab. 13. Two quantitive examples for image-based ReID are visualized by Fig. 7.
Figure 7: Rank-50 Retrieval Example for Image-Based Vehicle ReID using the SBS [14] Model Trained on the RoundaboutHD Subset. The blue bounding box marks the query image, green bounding boxes indicate correctly retrieved matches (true positives), and red bounding boxes denote incorrect matches (false negatives).
Algorithm mAP Rank-1 Rank-5 mINP
BoT[24] 3.08 15.13 25.71 0.42
AGW[37] 2.71 8.94 16.68 0.60
SBS[14] 19.51 44.80 58.73 3.56
Table 12: Vehicle ReID performance. This table shows the results using models pre-trained only on ImageNet; all values are percentages.
Algorithm mAP Rank-1 Rank-5 mINP
BoT[24] 98.45 98.73 99.05 98.41
AGW[37] 97.86 97.76 99.05 93.74
SBS[14] 99.19 99.66 99.66 98.57
Table 13: Pre-trained Vehicle ReID performance. This table shows the performance of models trained on our custom dataset, all values are in percentage.
As shown in Tab. 12, all three models initially achieve low mAP when using only ImageNet-pretrained weights, highlighting the domain gap between generic image datasets and real-world vehicle ReID tasks. However, fine-tuning on the RoundaboutHD dataset leads to substantial performance gains across all models. As shown in Fig. 7, the trained model achieves near-perfect accuracy in matching vehicle instances captured from multiple viewpoints. Notably, most false negatives are caused by the limited number of query images per identity rather than inherent model deficiencies. This significant improvement can be largely attributed to the roundabout scenario, which naturally captures vehicles from diverse angles—even within the same camera—thus enriching the visual appearance space and facilitating more robust feature learning for image-based ReID tasks.
4.4 Multi-Camera Tracking
Multi-camera tracking (MCT) is the task for associating vehicle identities across different camera views. In most of studies, MCT builds upon the outputs of SCT, associating identities by comparing distances between deep embedding features. Some approaches also integrate camera topology and spatiotemporal reasoning to further constrain the search space, enabling more accurate reconstruction of complete cross-camera trajectories.
Following the release of CityFlow and the CVPR AI City Challenge workshop, there are several research focused on the MCVT problem and achieved good results. Notable examples include ELECTRICITY [27], the winning method of the 2020 AI City Challenge, Liu et al. [19], winner in 2021, and Yang et al. [36], winner in 2022. In our benchmark evaluation, we choose ELECTRICITY as the baseline method due to its general and reproducible pipeline. Unlike Liu and Yang’s approaches, which rely on manually annotated spatial-temporal zone-based information between cameras that are only available for the CityFlow dataset, ELECTRICITY can be more broadly applied to other datasets without requiring prior human annotations. During the evaluation on RoundaboutHD, we filter out all static vehicle trajectories to remove the influence of parked vehicles visible in the scene. For the cross-camera matching stage, we fine-tune the threshold distance to 12 and set a hard removal distance of 80. The performance of ELECTRICITY across multiple datasets is summarized in Tab. 14.
Dataset IDF1 IDP IDR
CityFlow 46.16 / /
Synthehicle 41.50 48.0 37.5
RoundaboutHD (Ours) 28.14 26.45 30.06
Table 14: ELECTRICITY Performance Evaluation on Different Datasets.
This result provides a baseline of how SOTA algorithm performances in different MCVT dataset, with IDF1 score serving as the primary metric. Moreover, it also proves the increased challenge posed by RoundaboutHD compared to existing dataset, which is largely attributed to nonlinear vehicle trajectories and frequent occlusions inherent in roundabout environments.
5 Conclusion
This paper presents RoundaboutHD, a comprehensive, high-resolution real-world dataset designed to advance research in MCVT. Unlike previous open-source datasets, RoundaboutHD addresses the lack of high-quality, practically representative data in this field. It introduces additional challenges such as nonlinear vehicle motion and frequent occlusions, which arise naturally from the roundabout scenario.
The dataset comprises video footage from four non-overlapping cameras, each offering 10 minutes of fully annotated 4K, 15 fps video. In total, it includes 549,909 annotated bounding boxes across 512 unique vehicle identities. In addition to the full MCVT dataset, we provide dedicated subsets for object detection, single-camera tracking, and image-based vehicle ReID. A unified evaluation framework is also released, along with baseline results using several SOTA algorithms across all tasks. Experimental results prove that RoundaboutHD presents realistic and substantial challenges, making it a valuable benchmark for the development and evaluation of future MCVT methods.
While RoundaboutHD provides high-quality and richly annotated data, its focus on a single scene may limit scenario diversity, which could affect model generalizability in more varied environments. To address this, future extensions will aim to incorporate a wider variety of conditions, including different lighting, weather, and traffic patterns. We hope that the release of RoundaboutHD will inspire further advancements in the MTMCT research community.